# LoRA-Contextualizing Adaptation of Large Multimodal Models for Long Document Understanding

### v1

@@ Appendix B Evaluation Metrics
We report generalized accuracy (G-Acc) from MMLongBench-Doc (Ma et al., 2024c), a GPT-dependent, rule-based evaluation protocol . Model responses are simplified using GPT-4o and scored based on answer-type-specific rules. However, G-Acc has two limitations: it introduces randomness from GPT’s stochastic outputs and relies on answer-type annotations, limiting its applicability across datasets.
@@ Appendix C Example of Inference Failure Scenario
Figure C.1: Inference example of a challenging case in the SlideVQA dataset. LoCAL-Paligemma retrieved the wrong evidence page due to limitations in its retrieval module, leading to an incorrect answer. LoCAL-Phi-3-V retrieved the correct page but provided a wrong answer due to limitations in its QA module. Meanwhile, LoCAL-InternVL2-4B also assigned the highest relevance score to an incorrect page. However, since it processes multiple pages (top 5), the correct evidence page was included in the input, allowing its fine-tuned QA module to deliver the correct answer.

### v2

@@ Appendix B Evaluation Metrics
We report generalized accuracy (G-Acc) from MMLongBench-Doc (Ma et al., 2024d), a GPT-dependent, rule-based evaluation protocol . Model responses are simplified using GPT-4o and scored based on answer-type-specific rules. However, G-Acc has two limitations: it introduces randomness from GPT’s stochastic outputs and relies on answer-type annotations, limiting its applicability across datasets.
@@ Appendix C Example of Inference Failure Scenario
Figure C.1: Inference example of a challenging case in the SlideVQA dataset. SV-RAG-Paligemma retrieved the wrong evidence page due to limitations in its retrieval module, leading to an incorrect answer. SV-RAG-Phi-3-V retrieved the correct page but provided a wrong answer due to limitations in its QA module. Meanwhile, SV-RAG-InternVL2-4B also assigned the highest relevance score to an incorrect page. However, since it processes multiple pages (top 5), the correct evidence page was included in the input, allowing its fine-tuned QA module to deliver the correct answer.
@@ Appendix D Qaulitative Results in Question-Answering
Appendix E Examples from the VisR-Bench dataset
Figure E.1: Example question-and-answer pair from the VisR-Bench dataset, highlighting the reliance on both image and surrounding text for accurate responses.
Figure E.2: Example question-and-answer pair from the VisR-Bench dataset, highlighting the reliance on both image and surrounding text for accurate responses.
Appendix F Comparison of Retrieval Method Efficiency
Text Extraction Text Encoding Multimodal Encoding
PaddleOCR 0.275 BM25 0.0001 CLIP 0.022
BGE-m3 0.131 SigLip 0.109
PDF Parser 0.762 BGE-large 0.137 Col-Paligemma 0.140
NV-embed-v2 0.117 Col-Phi-3-V 0.230
Col-InternVL2 0.581
Table F.1: Per-page time cost of retrieval methods: The left table presents time cost (seconds) of text-based methods that rely on text extraction techniques, such as OCR models, followed by text encoders to compute page embeddings. The right table presents time cost (seconds) of multi-modal methods that encode the entire page as an image.
Appendix G Additional Experiment results
We compare our method with text-only baselines using a document parser [^3] Adobe Extract API: https://developer.adobe.com/document-services/apis/pdf-extract/ to highlight the advantages of MLLMs in multi-modal understanding. QA results are reported for the VisR-Bench and MMLongBench-Doc datasets, where PDF files are available.
Table G.1 presents QA results on VisR-Bench and MMLongBench-Doc datasets. To evaluate answer quality for VisR-Bench, where true answers are long and detailed, we introduce the Mean GPT Score (MGS), as string-matching methods often penalize variations in wording for lengthy answers. Instead, we prompt GPT-4o to compare a model’s answer with the ground truth and assign a binary score based on detail alignment.
QA Module Retrieval Module Evidence VisR-B MMLong
MGS G-Acc
Text only QA methods
Phi-3 + parser Col-Phi-3-V R5 14.1 29.2
GPT-4o + parser Col-Phi-3-V R5 24.9 43.2
GPT-4o + parser - A 27.6 42.4
MLLM QA models
PaliGemma Col-PaliGemma R1 12.2 23.9
Phi-3-V Col-Phi-3-V R1 24.2 30.7
SV-RAG-InternVL2 Col-InternVL2 R5 25.2 33.2
GPT-4o Col-Phi-3-V R5 47.2 55.1
GPT-4o - A 43.2 54.5
Table G.1: parser results
Our results indicate that using image evidence consistently outperforms text-only evidence. On VisR-Bench, text-only baselines showed a significant performance drop, emphasizing the dependency of questions on both image and text. However, the MGS of text-only baselines is not zero, likely because the model leverages text from a broader context rather than relying solely on the surrounding text, enabling it to extract relevant information even in the absence of visual input.
Additionally, reducing input pages with the retrieval module improved GPT-4o’s performance with image evidence, aligning with the findings in Table 2. In contrast, retrieval did not enhance GPT-4o’s performance on VisR-Bench in the text-only setting, likely because the evidence pages lacked sufficient information to fully address the questions. Including additional context in such cases might yield better results.